# A new approach to constraining properties of AGN host galaxies by combining image and SED decomposition: testing upon the relation

### v1

@@ 4.1 Stellar mass estimation
where refers to the difference between the observed black hole mass and the one inferred from the stellar mass with the local relation (Kormendy & Ho, 2013), hereafter KH13 relation, while is the evolution factor. We define to be the intrinsic scatter in BH mass. In the sampling process, we generated 100,000 mock AGNs for each set of (, ) values. We assigned those AGNs with uniformly distributed redshift in the range of , then randomly sampled the black hole mass of each AGN () according to the BHMF. Based on the given redshift and black hole mass of each mock AGN, we then sampled the Eddington ratio to obtain their bolometric luminosity. We removed the mock AGNs which have lower bolometric luminosity than which was the lower limit of our sample, in order to make the mock sample resemble the observed sample as much as possible. Then we assigned the stellar mass () of each mock sample based on the black hole masses and the intrinsic scatter in BH mass which was added as a random normal variable in the model.
@@
Based on the likelihood defined above, we derived the posterior distribution, as shown in Figure 5, through MCMC inference assuming flat prior. Note that some similar studies acquired divergent results under flat prior, thereby they obtained their final result through assuming a log-normal prior on in order to constrain the parameter space (Ding et al., 2020; Li et al., 2021a). Compared with those studies, we derive much higher intrinsic scatter of (former ), which may indicate that the black hole and the host galaxy do not grow in lockstep out to (Yang et al., 2018).
The relation is plotted in Figure 6. We adopted medians from the Bayesian SED stellar mass estimation, and the error bars represent the interval between the 16th and 84th percentiles of the Bayesian inference sample. We reconstructed a mock local sample with and and then applied the same luminosity cut as our observed sample to represent the distribution of the mock sample assuming that the correlation does not evolve at all. Using the model with and , which is the aforementioned best inference parameter set, the distribution of the simulated data points shown as orange contours is consistent with the observed ones. The orange contours span wider range compared with the blue ones, which is maily due to the larger intrinsic scatter. We note that there are 7 “BH monsters” lying above the red dashed line, with . If we ignore these “BH monsters”, the other 17 sources lie evenly around KH13 relation with median difference in of 0.43 dex.
@@ 5.1 Possible sources of error in our work
Firstly, the multiband fluxes which we use in the SED fitting routine are retrieved from various archival catalogs, which may suffer from inconsistency in photometry. For example, we adopt optical fluxes from Hubble Souce Catalog (Whitmore et al., 2016) measured with fixed aperture photometry. In the similar optical bands, however, SDSS catalogs only provide fluxes based on modelMag photometry for extended sources, which were derived based on model fitting of De vaucouleur profile or exponential profile. Such inconsistencies may account for the observed fluctuations in the data points across the optical bands of the SED (Figure 3). Data from ground-based infrared telescopes such as VISTA and UKIRT suffer from low resolution and insufficient pointing precision, which is likely to cause the fluxes to include the contamination from nearby objects.
In the AGN-host galaxy image decomposition procedure, the crucial part is PSF selection (Ding et al., 2020; Tanaka et al., 2024). Mistaking extended sources for PSFs could significantly bias the flux decomposition results, particularly given the limited quality of the imaging data used in our work. Despite this challenge, we employ a weighted method similar to that proposed by Ding et al. (2020), which utilizes multiple PSFs to mitigate the impact of potential mismatched PSFs. Note that Ding et al. (2020) collected all available star-like sources around their 32 sources to construct a PSF library. For each source, they conducted image decomposition using every PSF in the library, ranking the resulting and selecting the top N results in order to cover all possible PSF cases. However, in five out of ten bands which we used for image decomposition the number of available images is less than 5, posing challenges in building comprehensive PSF libraries for these bands. In order to maintain consistency throughout our investigation, we employ an in-between approach for image decomposition, selecting PSFs only from the neighboring regions for each target. While this approach may not capture the full variability in PSFs, it enables us to represent the relative goodness of image decomposition for different sources, which is preferable to coarsely fixing all uncertainties at 10%. Nevertheless, despite setting a lower limit on the estimation error, mismatched PSFs, if present, could still bias the fluxes of the stellar component towards lower values.
@@ 5.2 Constraining stellar masses
The primary advancement in our method compared to previous multi-band SED decomposition approaches (Zou et al., 2019; Suh et al., 2019, 2020) is the use of image decomposition results as additional constraints to refine SED models. We examined the posterior distribution of stellar masses with and without these extra constraints, denoted as orange and blue curves in Figure 7. Compared to the scenario without constraints, the posterior probability distribution functions (pdfs) with constraints span narrower wings, indicating that the stellar mass is better constrained.
@@
Although image decomposition could pose strong constraints upon host properties, obtaining high-resolution imaging data for high redshift samples often requires significant observing time with space telescopes, limiting the sample size in studies on high-redshift AGN host galaxies. On the other hand, broadband photometric data can be relatively more accessible. Therefore, we further assessed the relative constraints of broadband photometry in different bands. Given the considerable uncertainty in FIR observations, we experimented with a two-component model when fitting the data, excluding Herschel Pacs and SPIRE FIR data from the input. The corresponding pdfs are shown as green curves in Figure 7. For CID 71, excluding FIR photometry significantly enlarges the estimation error. However, for ECDFS 379, there is only one photometric point in the FIR bands, so excluding FIR makes no difference in the results. As detailed in Section 3.2, the cold dust model is based on the energy balance assumption, thus FIR data provides supplement information about the stellar population. We emphasize that the MIR data offers crucial details about the torus (Netzer, 2015) and can put strong constraints on the AGN component, which may help in mitigating the degeneracy between AGN and the stellar component in optical bands. The optical and NIR fluxes can meanwhile provide supplemental information when the image-decomposed bands are insufficient.
@@
Considering that throughout our study we find no clear bias between our sample with 1 or 2 bands of imaging and those with 3 or more bands, we infer that 2 or 3 bands of imaging is sufficient. We propose that future works could base on image decomposition for high-quality imaging data in as few as 3 optical-NIR bands and broadband photometry in other bands, thus expanding the sample size to increase statistical significance. With JWST surveys in NIR-MIR (), investigations on type-I AGNs with much larger samples would become possible.
@@ 6 Summary
In this work, we proposed a novel method combining image decomposition and SED decomposition to constrain stellar masses. This method enables us to take advantage of limited high-resolution imaging data and mitigate the degeneracy between AGN bright UV-optical emission and host stellar emission as much as possible, constraining host properties in multiple ways. To further test the viability of this method, we obtained a sample of 24 X-ray selected type-I AGNs, with redshift ranging from 0.73 to 2.47.
The main findings of this work are listed below:
We compared our stellar mass estimates with previous works which have samples overlapping with ours. We find that our results partially agree with those based on broadband SED decomposition (Zou et al., 2019; Suh et al., 2020), and are generally consistent with those based on image decomposition (Ding et al., 2020).
We studied the relation of our sample, finding much higher intrinsic scatter for our sample compared to the local relation, which may be caused by 7 “BH monsters” in our sample. If these “BH monsters” are excluded, our results basically agree with the local correlation given in Kormendy & Ho (2013).
We evaluated the impact of various data on the derived posterior pdf of stellar masses. The results show that stellar mass is better constrained when using our method which combines image decomposition data and MIR/FIR broadband photometric data.
@@ Appendix A Image decomposition results
Note. — (a) “” means that data is not available for this band. (b) “” means that this flux is an upper limit.
@@
Note. — “” means that data is not available for this band.

### v2

@@ 4.1 Stellar mass estimation
where refers to the difference between the observed black hole mass and the one inferred from the stellar mass with the local relation (Kormendy & Ho, 2013), hereafter KH13 relation, while is the evolution factor. We define as the intrinsic scatter in the BH mass. In the sampling process, we generated 100,000 mock AGNs for each set of (, ) values. We assigned those AGNs with uniformly distributed redshift in the range of , then randomly sampled the black hole mass of each AGN () according to the BHMF. Based on the given redshift and black hole mass of each mock AGN, we then sampled the Eddington ratio to obtain their bolometric luminosity. We removed mock AGNs that have lower bolometric luminosity than which was the lower limit of our sample, to make the mock sample resemble the observed sample as much as possible. Then we assigned the stellar mass () of each mock sample based on the black hole masses and the intrinsic scatter in BH mass , which was added as a random normal variable in the model.
@@
Based on the likelihood defined above, we derived the posterior distribution, as shown in Figure 5, through MCMC inference assuming a flat prior. Note that some similar studies acquired divergent results under flat prior; thus, they obtained their final result by assuming a log-normal prior on in order to constrain the parameter space (Ding etĀ al., 2020; Li etĀ al., 2021a). Compared with those studies, we derive much higher intrinsic scatter of (formerly ), which may indicate that the black hole and the host galaxy do not grow in lockstep with (Yang etĀ al., 2018).
The relation is plotted in Figure 6. We adopted medians from the Bayesian SED stellar mass estimation, and the error bars represent the interval between the 16th and 84th percentiles of the Bayesian inference sample. We reconstructed a mock local sample with and and then applied the same luminosity cut as our observed sample to represent the distribution of the mock sample assuming that the correlation does not evolve at all. Using the model with and , which is the set of aforementioned best inference parameters, the distribution of the simulated data points shown as orange contours is consistent with the observed ones. The orange contours span a wider range compared with the blue ones, which is most likely due to the larger intrinsic scatter. We note that there are 7 āBH monstersā lying above the red dashed line, with . If we ignore these āBH monstersā, the other 17 sources lie evenly around the KH13 relation with a median difference in of 0.43 dex.
@@ 5.1 Possible sources of error in our work
Firstly, the multiband fluxes that we use in the SED fitting routine are retrieved from various archival catalogs, which may suffer from inconsistency in photometry. For example, we adopt optical fluxes from Hubble Souce Catalog (Whitmore etĀ al., 2016) measured with fixed aperture photometry. However, in similar optical bands, SDSS catalogs only provide fluxes based on modelMag photometry for extended sources, which were derived based on model fitting of the De vaucouleur profile or exponential profile. Such inconsistencies may explain the observed fluctuations in data points in the optical bands of the SED (Figure 3). Data from ground-based infrared telescopes such as VISTA and UKIRT suffer from low resolution and insufficient pointing precision, which is likely to cause the fluxes to include contamination from nearby objects.
In the AGN-host galaxy image decomposition procedure, the crucial part is PSF selection (Ding etĀ al., 2020; Tanaka etĀ al., 2024). The mistake of using extended sources for PSFs could significantly bias the flux decomposition results, particularly given the limited quality of the imaging data used in our work. Despite this challenge, we employ a weighted method similar to that proposed by Ding etĀ al. (2020), which utilizes multiple PSFs to mitigate the impact of potential mismatched PSFs. Note that Ding etĀ al. (2020) collected all available star-like sources around their 32 sources to construct a PSF library. For each source, they performed an image decomposition using every PSF in the library, ranking the resulting and selecting the top N results to cover all possible PSF cases. However, in five out of ten bands that we used for image decomposition, the number of available images is less than 5, posing challenges in building comprehensive PSF libraries for these bands. In order to maintain consistency throughout our investigation, we employ an in-between approach for image decomposition, selecting PSFs only from the neighboring regions for each target. Although this approach may not capture the full variability in PSFs, it enables us to represent the relative goodness of image decomposition for different sources, which is preferable to coarsely fixing all uncertainties at 10%. Nevertheless, despite setting a lower limit on the estimation error, mismatched PSFs, if present, could still bias the fluxes of the stellar component towards lower values.
@@ 5.2 Constraining stellar masses
The primary advancement in our method compared to previous multiband SED decomposition approaches (Zou etĀ al., 2019; Suh etĀ al., 2019, 2020) is the use of image decomposition results as additional constraints to refine SED models. We examined the posterior distribution of stellar masses with and without these extra constraints, denoted as orange and blue curves in Figure 7.
@@
Although image decomposition could pose strong constraints upon host properties, obtaining high-resolution imaging data for high-redshift samples often requires significant observing time with space telescopes, limiting the sample size in studies on high-redshift AGN host galaxies. On the other hand, broadband photometric data can be relatively more accessible. Therefore, we further assessed the relative constraints of broadband photometry in different bands. Given the considerable uncertainty in FIR observations, we experimented with a two-component model when fitting the data, excluding Herschel Pacs and SPIRE FIR data from the input. The corresponding pdfs are shown as green curves in Figure 7. For CID 71, excluding FIR photometry significantly increases the estimation error. However, for ECDFS 379, there is only one photometric point in the FIR bands, so excluding FIR does not make any difference in the results. As detailed in Section 3.2, the cold dust model is based on the energy balance assumption, thus FIR data provides supplement information about the stellar population. We emphasize that the MIR data provide crucial details about the torus (Netzer, 2015) and can put strong constraints on the AGN component, which may help in mitigating the degeneracy between AGN and the stellar component in optical bands. The optical and NIR fluxes can meanwhile provide supplementary information when the image-decomposed bands are insufficient.
@@
Considering that throughout our study we find no clear bias between our sample with 1 or 2 bands of imaging and those with 3 or more bands, we infer that 2 or 3 bands of imaging is sufficient. We propose that future work could be based on image decomposition for high-quality imaging data in as few as 3 optical-NIR bands and broadband photometry in other bands, thus expanding the sample size to increase statistical significance. With JWST surveys in NIR-MIR (), investigations on type I AGNs with much larger samples would become possible.
5.3 SFR diagnostics
We further tested the star formation rates (SFR) obtained from our SED fitting results using the parameter, which quantifies the āstarburstinessā of an AGN compared to main-sequence (MS) star-forming galaxies. This parameter is defined as the ratio of the SFR of an AGN and that of MS galaxies of similar and : . Following Cristello etĀ al. (2024), who investigated the dependence of on and using CIGALE SED fitting results for AGNs (Zou etĀ al., 2022), we analyze the SFR for our samples.
Considering that FIR data are crucial for accurately constraining SFR, samples without FIR photometry were excluded in this analysis. We used the MS data provided by Cristello etĀ al. (2024), which was derived based on a specific star formation rate (sSFR) threshold method. Specifically, we adopted the MS at to approximately match the redshift distribution of our samples. Since their SFR estimates are derived using similar SED fitting procedures, the adopted MS should be consistent with our sample. In this subsection, we employ a delayed SFH and a non-evolving chemical abundance model to estimate SFR, which is consistent with Zou etĀ al. (2022). Figure 8 presents the of our sample, together with the relation for high luminosity () AGNs provided by Cristello etĀ al. (2024). This range roughly corresponds to the luminosity of our presented sample, which spans from 43.5 to 44.5. We found that host galaxies with tend to exhibit higher SFR than their MS counterparts, whereas more massive galaxies with show a mean approximately equal to unity. This trend, where less massive AGN hosts tend to exhibit more enhanced star formation, has also been reported by Cristello etĀ al. (2024). However, our data show a larger scatter compared with the literature, which could be due to the limited sample size. It is worth noting that the delayed SFH was applied only for the SFR estimation, while a double-power law SFH was adopted throughout the rest of this work. When comparing these two SFH models, we found no systemic bias in the derived stellar masses. However, the double power law SFH provided better agreement with other works in Section 4.1.
@@ 6 Summary
In this work, we proposed a novel method that combines image decomposition and SED decomposition to constrain stellar masses. This method enables us to take advantage of limited high-resolution imaging data and mitigate the degeneracy between AGN bright UV-optical emission and host stellar emission as much as possible, constraining host properties in multiple ways. To further test the viability of this method, we obtained a sample of 24 type I AGNs selected by X-ray observations, with redshift ranging from 0.73 to 2.47.
The main findings of this work are listed below.
ā¢
We compared our stellar mass estimates with previous works which have samples overlapping with ours. We find that our results are partially in agreement with those based on the broadband SED decomposition (Zou etĀ al., 2019; Suh etĀ al., 2020) and are generally consistent with those based on the image decomposition (Ding etĀ al., 2020).
ā¢
We studied the relation of our sample, finding much higher intrinsic scatter for our sample compared to the local relation, which may be caused by 7 āBH monstersā in our sample. If these āBH monstersā are excluded, our results basically agree with the local correlation given in Kormendy & Ho (2013).
ā¢
We evaluated the impact of various data on the derived posterior pdf of stellar masses. The results show that stellar mass is better constrained when using our method, which combines image decomposition data and MIR/FIR broadband photometric data.
ā¢
We estimated the SFR for samples with FIR photometry and found that less massive AGN hosts () tend to exhibit higher SFR compared to MS galaxies, which is consistent with recent studies.
@@ Appendix A Image decomposition results
Note. ā (a) āā means that data is not available for this band. (b) āā means that this flux is an upper limit.
@@
Note. ā āā means that data is not available for this band.